OpenStudy (abdela25):

HELP PLEASE.... MEDALS WILL BE REWARDED TO FIRST AND BEST ANSWER...small math word problem

OpenStudy (abdela25):

This is the word problem the debate club needs $240.00 to attend a debate tournament the club decides to sell cups of iced tea and lemonade at baseball games. iced tea will be sold for $0.50 and lemonade will be sold for $0.80 per cup

OpenStudy (abdela25):

the first question to it is write an equation to find how many cups of each bevergae must be sold to rais $240.00 My answer to this one was: x=number of cups of iced tea y=number of cups of lemonade 0.5x+0.8y=240, 5x+8y+2400, 8y+-5x+2400

OpenStudy (abdela25):

the second question to this problem was: Graph the equation. what are the x- and y-intercepts my answer was: x-intercept- y=0, -5/8x+300=0, -5/8x=-300 x=-300/(-5/8), x=-300(-8/5), x=2400/5 x=480, x-intercept = (480,0) y-intercept- x=0, y=-5/8(0)+300, y=300 y-intercept = (0,300)

OpenStudy (abdela25):

but this part of the question also acts for a graph i dont know how to graph this

OpenStudy (anonymous):

5

OpenStudy (anonymous):

jk

OpenStudy (anonymous):

The first equation is almost right: \(0.5x+0.8y=240\) This would probably be an acceptable answer. \(5x+8y=2400\) I assume this was what you meant for the second step and would also be acceptable. \(8y+-5x+2400 \) I may be confused on the question, but this doesn't make any sense to me. I think you should have stopped at step two.

OpenStudy (anonymous):

To graph the equation, first solve for y.

OpenStudy (anonymous):

And thank you for showing your attempt when asking the question. That shows honest effort.

OpenStudy (abdela25):

8y+-5x=2400 i mean to put an = sign and not a + i did this to put 8y first

OpenStudy (anonymous):

I don't understand why you are subtracting \(5y\).

OpenStudy (abdela25):

and i think i did solve for y in the second part of the question which was y=-5/8x+300

OpenStudy (anonymous):

\(5x+8y=8y+5x\)

OpenStudy (abdela25):

so was my solving for y correct

OpenStudy (anonymous):

Still checking, it's kind of hard to read in paragraph form.

OpenStudy (abdela25):

oh sorry about that

OpenStudy (anonymous):

\(5x+8y=2400\) \(8y=-5y+2400\) \(y=-\frac{5}{8}y+300\) Yes, you solved for y correctly. Now I also see where the -5x comes in. Sorry.

OpenStudy (anonymous):

The y's on the right should be x's.

OpenStudy (abdela25):

ok thanks and yeah sorry about the paragraph typing

OpenStudy (anonymous):

Live and learn. No problem.

OpenStudy (abdela25):

so how would i graph this equation and also the second question for this problem asks to graph and find x- and y-intercepts i think i did that but i cant graph it

OpenStudy (anonymous):

\(0=-\frac{5}{8}x+300\) \(-\frac{5}{8}x=300\) \(x=-480\) y=intercept - also correct. \(5x+8y=2400\) \(8y=2400\) \(y=300\) x-intercept - positive, not negative.

OpenStudy (anonymous):

Are those the values that you had?

OpenStudy (abdela25):

so i think thats what i put correct? or was three any problems with it ?

OpenStudy (anonymous):

I think you may have dropped the negative when you typed it in. Not a problem as long as it matches what is on your paper in front of you.

OpenStudy (anonymous):

Or I added too many negatives. (480,0) and (0,300) are the intercepts.

OpenStudy (abdela25):

ok thanks so do i graph using the equation found in question 1 or in question 2 there is also question 3 for this problem but to answer it i need the graph from question 2

OpenStudy (anonymous):

As for the graph.... \(y=-\frac{5}{8}y+300\) Start with the y-intercept at (0,300). The slope is \(-\frac{5}{8}\) or \(\frac{rise}{run}\) You have a negative so count up 5 units and left 8 units to find the next point. Left for a negative slope and right for a positive slope.

OpenStudy (abdela25):

small question before we continue in your equation you have y= -5/8y+300 shouldnt it be y= -5/8x +300

OpenStudy (anonymous):

Yes it should.

OpenStudy (anonymous):

Thank you for noticing the error. That shows you are thinking on your own and learning!

OpenStudy (abdela25):

your welcome and yes i do want to learn as much as possible so i can be able to do it on my own in the future

OpenStudy (abdela25):

but graphing for me was always hard and still is hard thats one thing in math that i cannot understand yet

OpenStudy (anonymous):

\(y=mx+b\) m is the slope and b is the y-intercept

OpenStudy (anonymous):

I would suggest drawing the graph in blocks of 100 units.

OpenStudy (anonymous):

Start with the point (300,0) the y-intercept.